OpenStudy (callisto):

Sequence question#2 - a childish one

OpenStudy (callisto):

I can work out the answer. The only thing I would like to see is your workings.

OpenStudy (unklerhaukus):

A. P. means: arithmetic progression?

OpenStudy (lgbasallote):

tsk childish she says -____- haha

OpenStudy (callisto):

yes

OpenStudy (callisto):

I did it like this: Sum of the 5 terms = (3+23)*5/2 = 65 a+b+c = 65 - 3-23 = 39 Any faster method?

OpenStudy (unklerhaukus):

by inspection \[\{3,a,b,c,23\}=\{3,a,13,c,23\}=\{3,8,13,18,23\}\] \[a+b+c=8+13+18=39\]

OpenStudy (experimentx):

a+b+c = 3b b = (3+23)/2

OpenStudy (unklerhaukus):

b has to be halfway in-between 3, and 23. so b =13

OpenStudy (callisto):

@experimentX Wow~ another cute solution!!! @UnkleRhaukus Thanks!~

OpenStudy (experimentx):

@UnkleRhaukus you are a physicist ... aren't you??

OpenStudy (unklerhaukus):

i am mid way through my bachelor of science (nuclear science technologies) degree

OpenStudy (experimentx):

thats cool man ...!!!

OpenStudy (unklerhaukus):

thankyou very much, what stage in your career are you at?

OpenStudy (anonymous):